# MedCare: Advancing Medical LLMs through Decoupling Clinical Alignment and Knowledge Aggregation

### v2

@@ Abstract
Large language models (LLMs) have shown substantial progress in natural language understanding and generation, proving valuable especially in the medical field. Despite advancements, challenges persist due to the complexity and diversity inherent in medical tasks, which can be categorized as knowledge-intensive tasks and alignment-required tasks. Previous approaches either ignore the latter task or focus on a minority of tasks and hence lose generalization. To address these drawbacks, we propose a progressive fine-tuning pipeline. This pipeline employs a Knowledge Aggregator and a Noise Aggregator to encode diverse knowledge in the first stage and filter out detrimental information. In the second stage, we drop the Noise Aggregator to avoid the interference of suboptimal representation and leverage an additional alignment module optimized towards an orthogonal direction to the knowledge space to mitigate knowledge forgetting. Based on this two-stage paradigm, we proposed a Medical LLM through decoupling Clinical Alignment and Knowledge Aggregation (MedCare), which is designed to achieve state-of-the-art (SOTA) performance on over 20 medical tasks, as well as SOTA results on specific medical alignment tasks. Various model sizes of MedCare (1.8B, 7B, 14B) all demonstrate significant improvements over existing models with similar model sizes.
@@
Yusheng Liaothanks: Equal contribution.,1, Shuyang Jiang∗,2,3, Yanfeng Wang1,3, Yu Wangthanks: Corresponding Author,1,3 1Shanghai Jiao Tong University 2Fudan University 2Shanghai Artificial Intelligence Laboratory {liao20160907,,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn shuyangjiang23@m.fudan.edu.cn
@@ 1 Introduction
Although progress has been made, LLMs still face challenges due to the complexity and diversity of medical tasks. This is particularly evident when data is intentionally structured to represent a diverse set of tasks (Chung et al., 2022; Wei et al., 2021; Sanh et al., 2021). We categorize medical tasks into two types: knowledge-intensive tasks and alignment-required tasks. Knowledge-intensive tasks, such as medical question answering (Zhang et al., 2018) and medical dialogues (Liu et al., 2022a; Zhao et al., 2022), primarily test the model’s understanding of internal medical knowledge. Conversely, alignment-required tasks, like clinical terminology standardization (Zhang et al., 2022b) and medical entity recognition (Hongying et al., 2021), demand not only medical knowledge but also strict adherence to output formats. Figure 1 illustrates these two categories.
However, previous medical LLMs, such as HuatuoGPT (Wang et al., 2023a), HuatuoGPT-II (Chen et al., 2023a), Med-PaLM (Singhal et al., 2023a) and BioMistral (Labrak et al., 2024), have primarily focused on enhancing the encoding of medical knowledge in LLMs, neglecting the requirements of alignment-required tasks Van Veen et al. (2024). On the other hand, alignment-oriented methods Liu et al. (2023b); Cai et al. (2023); Wang et al. (2023a) that perform alignment through fine-tuning often suffer from an "alignment tax". This results in knowledge forgetting and performance drops in knowledge-intensive tasks (Lin et al., 2023; Gekhman et al., 2024). These issues limit the practical use of LLMs in healthcare.
@@ 2 Preliminaries
$\mathbf{H}=W\mathbf{X}+\mathbf{X}\Delta W=\mathbf{X}W+\frac{\alpha}{r}\mathbf{X}AB$ (1)
where $W\in\mathbb{R}^{d\times d^{\prime}},A\in\mathbb{R}^{d\times r},B\in\mathbb{R}^{r\times d^{\prime}}$ and $r\ll d,d^{\prime}$. $\alpha$ is the scaling factor. Without loss of generality, we omit the layer index for the following formula. At the beginning of training, $B$ is initialized to an all-zero matrix and $A$ uses Gaussian initialization to make sure that the product $AB$ is zero at initialization.
@@
$\mathbf{X}=(\mathbf{X}W_{u}\cdot\mathrm{silu}(\mathbf{X}W_{g}))W_{d}$ (2)
where $W_{g}\in\mathbb{R}^{d^{\prime}\times d},W_{u}\in\mathbb{R}^{d^{\prime}\times d},W_{d}\in\mathbb{R}^{d\times d^{\prime}}$ and $d^{\prime}=8/3d$. In the LoRA forward passes, each linear layer $W\in\{W_{d},W_{u},W_{g}\}$ is updated using the LoRA module described in Eq.1. Consider an MoLoRA module with $E$ experts, and each expert is denoted as $\{E_{i}\}_{i=1}^{E}$, the forward pass of the linear layer is formulated as:
@@
where $\mathbf{H}$ is the processed output and $G(\cdot)=\mathrm{softmax}(\mathbf{X}W_{r}),W_{r}\in\mathbb{R}^{d\times E}$ is the router in MoLoRA with top-$K$ selection which holds the top-$K$ affinity with respect to current input $\mathbf{X}$. Assuming each LoRA expert shares the same rank $r$ and $\alpha$ and we obtain the weight $w_{i}$ from the router $G(\cdot)$ as $w_{i}=G(\mathbf{X})_{i}$, the overall output of the MoLoRA linear layer can be formulated as:
@@ 3.2 Miscellaneous Knowledge Aggregation
$\mathbf{X}$ $=\mathbf{X}W+\mathrm{KA}(\mathbf{X})+\mathrm{NA}(\mathbf{X})$ (6)
@@
$\mathbf{X}=\mathbf{X}W+\mathrm{KA}(\mathbf{X})$ (8)
@@ 3.3 Downstream Alignment
$\mathbf{X}$ $=\mathbf{X}W+\mathrm{KA}(\mathbf{X})+\mathrm{Align}(\mathbf{X})$ (9)
@@ 5 Discussion
To demonstrate this, we tested the models with two parallel LoRA modules to investigate the role of the Noise Aggregator. As shown in Table 4, the parallel LoRA module failed to decouple the learning process due to its symmetric structure. Conversely, in the structure of MedCare, the Knowledge Aggregator predominantly acquires common knowledge from the fine-tuning corpus, while the Noise Aggregator focuses more on learning alignment formats.
@@ Appendix A Related Works
Full fine-tuning effectively adapts base large language models to downstream tasks but also consumes significant computational resources with the increasing size of models and the number of tasks. To address this, Parameter-Efficient Fine-Tuning (PEFT) methods have been introduced. These methods freeze the base language models, modifying only a negligible number of parameters during the training phase, yet achieving similar or even superior performance with limited fine-tuning data. Among these methods, Adapter-Tuning (Rebuffi et al., 2017; Houlsby et al., 2019; Lin et al., 2020; Pfeiffer et al., 2021) was the pioneering architecture that connected two additional projection layers to the pretrained language model. In addition to incorporating additional modules, Prefix-tuning (Li and Liang, 2021) introduces learnable prefix tokens and prepends them before the input prompts. Differently, Prompt-Tuning (Lester et al., 2021) utilizes learnable prompt tokens for each task, as a multi-task PEFT approach in NLU scenarios. Following these, P-Tuning (Liu et al., 2023c) and P-Tuning-v2 (Liu et al., 2022b) move away from explicit prompts and employ a prompt-generator to convert pseudo prompts into task prompts, allowing decoder-only models to also perform NLU tasks. However, these methods introduce additional priors and significant inference latency. In contrast, Low-Rank Adaptation (LoRA)(Hu et al., 2022) and its variant, Weight-Decomposed Low-Rank Adaptation (DoRA)(Liu et al., 2024b), take a different approach. LoRA updates original parameters with two low-rank matrices without assuming any specific task or architecture, eliminating inference latency by merging back these two matrices to the original weight. DoRA extends this by incorporating weight decomposition, achieving performance comparable to full fine-tuning. Nonetheless, transferring these methods to multi-task learning scenarios without manual adjustments remains a challenge.

### v3

@@ Abstract
Large language models (LLMs) have shown substantial progress in natural language understanding and generation, proving valuable especially in the medical field. Despite advancements, challenges persist due to the complexity and diversity inherent in medical tasks, which can be categorized as knowledge-intensive tasks and alignment-required tasks. Previous approaches either ignore the latter task or focus on a minority of tasks and hence lose generalization. To address these drawbacks, we propose a progressive fine-tuning pipeline. This pipeline employs a Knowledge Aggregator and a Noise Aggregator to encode diverse knowledge in the first stage and filter out detrimental information. In the second stage, we drop the Noise Aggregator to avoid the interference of suboptimal representation and leverage an additional alignment module optimized towards an orthogonal direction to the knowledge space to mitigate knowledge forgetting. Based on this two-stage paradigm, we proposed a Medical LLM through decoupling Clinical Alignment and Knowledge Aggregation (MedCare), which is designed to achieve promising performance on over 20 medical tasks, as well as results on specific medical alignment tasks. Various model sizes of MedCare (1.8B, 7B, 14B) all demonstrate significant improvements over existing models with similar model sizes. Our code and datasets are available at https://github.com/BlueZeros/MedCare.
@@
Yusheng Liaothanks: Equal contribution.,♠,♢, Shuyang Jiang∗,♣,♢, Zhe Chen♠,♢, Yanfeng Wang♠,♢, Yu Wangthanks: Corresponding Author,♠,♢ Shanghai Jiao Tong University Fudan University Shanghai Artificial Intelligence Laboratory {liao20160907,chenzhe2018,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn shuyangjiang23@m.fudan.edu.cn
@@ 1 Introduction
Although progress has been made, LLMs still face challenges due to the complexity and diversity of medical tasks. This is particularly evident when data is intentionally structured to represent a diverse set of tasks (Chung et al., 2022; Wei et al., 2021; Sanh et al., 2021). We categorize medical tasks into two types: knowledge-intensive tasks and alignment-required tasks. Knowledge-intensive tasks, such as medical question answering (Zhang et al., 2018) and medical dialogues (Liu et al., 2022a; Zhao et al., 2022), primarily test the model’s understanding of internal medical knowledge. Conversely, alignment-required tasks, like clinical terminology standardization (Zhang et al., 2022b) and medical entity recognition (Hongying et al., 2021), demand not only medical knowledge but also strict adherence to output formats. Figure 1 illustrates these two categories. In general scenarios, like medical question-answering or dialogue, the LLMs only need to answer the user’s questions correctly. However, tasks in clinical applications like report generation or situations usually require the model to produce a correct response and adhere to specific formats.
Previous medical LLMs, such as HuatuoGPT (Wang et al., 2023a), HuatuoGPT-II (Chen et al., 2023a), Med-PaLM (Singhal et al., 2023a) and BioMistral (Labrak et al., 2024), have primarily focused on enhancing the encoding of medical knowledge in LLMs, neglecting the requirements of alignment-required tasks Van Veen et al. (2024). On the other hand, alignment-oriented methods Liu et al. (2023b); Cai et al. (2023); Wang et al. (2023a) that perform alignment through fine-tuning often suffer from an "alignment tax". This results in knowledge forgetting and performance drops in knowledge-intensive tasks (Lin et al., 2023; Gekhman et al., 2024). These issues limit the practical use of LLMs in healthcare.
@@ 2 Preliminaries
$\mathbf{H}=\mathbf{X}W+\mathbf{X}\Delta W=\mathbf{X}W+\frac{\alpha}{r}\mathbf{X}AB$ (1)
where $W\in\mathbb{R}^{d\times d^{\prime}},A\in\mathbb{R}^{d\times r},B\in\mathbb{R}^{r\times d^{\prime}}$ and $r\ll d,d^{\prime}$. $\mathbf{H}$ is the processed output. $\alpha$ is the scaling factor. Without loss of generality, we omit the layer index for the following formula. At the beginning of training, $B$ is initialized to an all-zero matrix and $A$ uses Gaussian initialization to make sure that the product $AB$ is zero at initialization.
@@
$\mathbf{H}=(\mathbf{X}W_{u}\cdot\mathrm{silu}(\mathbf{X}W_{g}))W_{d}$ (2)
where $W_{g}\in\mathbb{R}^{d\times d^{\prime}},W_{u}\in\mathbb{R}^{d\times d^{\prime}},W_{d}\in\mathbb{R}^{d^{\prime}\times d}$, and $d^{\prime}=8/3d$. In the LoRA forward passes, each linear layer $W\in\{W_{d},W_{u},W_{g}\}$ is updated using the LoRA module described in Eq.1. Consider an MoLoRA module with $E$ experts, and each expert is denoted as $\{E_{i}\}_{i=1}^{E}$, the forward pass of the linear layer is formulated as:
@@
where $G(\cdot)=\mathrm{softmax}(\mathbf{X}W_{r}),W_{r}\in\mathbb{R}^{d\times E}$ is the router in MoLoRA with top-$K$ selection which holds the top-$K$ affinity with respect to current input $\mathbf{X}$. Assuming each LoRA expert shares the same rank $r$ and $\alpha$ and we obtain the weight $w_{i}$ from the router $G(\cdot)$ as $w_{i}=G(\mathbf{X})_{i}$, the overall output of the MoLoRA linear layer can be formulated as:
@@ 3.2 Miscellaneous Knowledge Aggregation
$\mathbf{H}$ $=\mathbf{X}W+\mathrm{KA}(\mathbf{X})+\mathrm{NA}(\mathbf{X})$ (6)
@@
$\mathbf{H}=\mathbf{X}W+\mathrm{KA}(\mathbf{X})$ (8)
@@ 3.3 Downstream Alignment
$\mathbf{H}$ $=\mathbf{X}W+\mathrm{KA}(\mathbf{X})+\mathrm{Align}(\mathbf{X})$ (9)
@@ 5 Discussion
  • RQ5: How does the effect of MedCare compare to other PEFT methods?
  • RQ6: How does MedCare perform in other languages?
@@
To demonstrate this, we tested the models with two parallel LoRA modules to investigate the role of the Noise Aggregator. As shown in Table 4, the parallel LoRA module failed to decouple the learning process due to its symmetric structure. However, it is surprising that removing one of the FFN LoRA modules can improve the model performance on knowledge-intensive tasks. Similar phenomena are also found in previous work (Jiang et al., 2024a). They found that removing the fine-tuned FFN parameters can fully utilize the general capacity of the base model by producing a more similar hidden state. In the structure of MedCare, the Knowledge Aggregator predominantly acquires common knowledge from the fine-tuning corpus, while the Noise Aggregator focuses more on learning alignment formats.
@@
Knowledge-Intensive Cross-Lingual Generalization
Model CMMLU CEval PLE Pha PLE TCM Avg. MedQA MedMCQA MMLU Avg.
Qwen1.5-7B 66.77 68.29 32.29 29.17 49.13 40.46 45.33 61.43 49.07
Huatuo-II-7B 59.08 62.40 47.70 47.50 54.17 40.69 44.75 44.33 43.26
LoRA (Hu et al., 2021) 65.21 60.98 46.67 50.83 55.92 40.53 44.94 61.24 48.90
LoRA+ (Hayou et al., 2024) 63.29 68.29 45.63 45.63 55.71 38.65 45.21 61.69 48.52
MoRA (Jiang et al., 2024b) 65.07 63.41 49.17 47.92 56.39 38.65 45.66 61.11 48.47
DoRA (Liu et al., 2024b) 65.14 70.73 50.00 50.83 59.18 39.75 46.02 60.60 48.79
MedCare 66.25 70.73 50.83 53.33 60.29 40.96 46.98 62.27 50.07
Table 5: Comparison with other parameter efficient fine-tuning (PEFT) methods on knowledge-intensive tasks and cross-lingual generalization. The results of PEFT methods are fine-tuned on Qwen1.5-7B. The datasets in cross-lingual generalization are all in English. Note that the ‘MedQA’ indicates the US subset.
@@
Response to RQ5: MedCare outperforms the commonly used PEFT methods on knowledge-intensive tasks.
We choose four PEFT methods to show the effectiveness of MedCare, including LoRA (Hu et al., 2021), LoRA+ (Hayou et al., 2024), MoRA (Jiang et al., 2024b), and DoRA (Liu et al., 2024b). It is observed in Table 5 that most of the PEFT methods suffer from performance drops on the medical parts of the general benchmarks CMMLU and CEval. This shows that domain adaptation will destroy the generalization ability of LLM and reduce its ability in scenarios outside the distribution of training data. The more effective methods, like DoRA and MedCare, can not only improve the domain capacity but also keep the generalization ability of the LLMs simultaneously, where MedCare demonstrates more superiority than the DoRA method.
Response to RQ6: MedCare shows strong cross-lingual generalization ability with only fine-tuned with monolingual data.
We choose three English medical datasets, MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and the medical part of MMLU (Hendrycks et al., 2021), to validate the cross-lingual capacity of the LLMs. The results in Table 5 show that all the other PEFT methods with only Chinese fine-tuning data achieve worse performance on English medical benchmarks when compared to the base model Qwen1.5-7B. However, with only Chinese fine-tuning data, MedCare can still improve the model’s performance on the English dataset with 1 point, showing strong cross-lingual ability and applicability.
@@ 6 Conclusions
Acknowledgements
This work is supported by the National Key R&D Program of China (No. 2022ZD0162101), the National Natural Science Foundation of China (No. 62106140), and STCSM (No. 21511101100, No. 22DZ2229005)
@@ Appendix A Related Works
Full fine-tuning effectively adapts base large language models to downstream tasks but also consumes significant computational resources with the increasing size of models and the number of tasks. To address this, Parameter-Efficient Fine-Tuning (PEFT) methods have been introduced. These methods freeze the base language models, modifying only a negligible number of parameters during the training phase, yet achieving similar or even superior performance with limited fine-tuning data. Among these methods, Adapter-Tuning (Rebuffi et al., 2017; Houlsby et al., 2019; Lin et al., 2020; Pfeiffer et al., 2021) was the pioneering architecture that connected two additional projection layers to the pretrained language model. In addition to incorporating additional modules, Prefix-tuning (Li and Liang, 2021) introduces learnable prefix tokens and prepends them before the input prompts. Differently, Prompt-Tuning (Lester et al., 2021) utilizes learnable prompt tokens for each task, as a multi-task PEFT approach in NLU scenarios. Following these, P-Tuning (Liu et al., 2023c) and P-Tuning-v2 (Liu et al., 2022b) move away from explicit prompts and employ a prompt-generator to convert pseudo prompts into task prompts, allowing decoder-only models to also perform NLU tasks. However, these methods introduce additional priors and significant inference latency. In contrast, Low-Rank Adaptation (LoRA)(Hu et al., 2022) and its variant, Weight-Decomposed Low-Rank Adaptation (DoRA)(Liu et al., 2024c), take a different approach. LoRA updates original parameters with two low-rank matrices without assuming any specific task or architecture, eliminating inference latency by merging back these two matrices to the original weight. DoRA extends this by incorporating weight decomposition, achieving performance comparable to full fine-tuning. Nonetheless, transferring these methods to multi-task learning scenarios without manual adjustments remains a challenge.